The 2012 financial statements for the Griffin Company are as follows:

                                    Griffin Company

                                    Statement of Financial Position

                                                            12/31/12                                  12/31/11

Assets

Cash

$40,000

$10,000

Accounts receivable

30,000

55,000

Inventory

110,000

70,000

Property, plant, and equipment

250,000

257,000

Total Assets

$430,000

$392,000

Limited and stockholders’ equity

Current liabilities

$40,000

$50,000

5% mortgage payable

30,000

162,000

Common Stock (30,000 shares)

110,000

150,000

Retained earnings

250,000

30,000

Total liabilities and stakeholder’s equity

$430,00

$392,000

            Griffin Company

Income statemen for the year ended December 31,2012

Sales on account

$420,000

Less expenses:

Cost of goods sold

$214,000

Salary expense

50,000

Depreciation Expense

7,000

Interest Expense

9,000

Total Expense

$280,000

Income before taxes

$140,000

Income tax expense (50%)

70,000

Net Income

$70,000

Compute the following ratios for the Griffin Company for the year ending December 31,2012:

  1. Profit margin ratio (before interest and taxes)
  2. Total asset turnover
  3. Rate of Return on total assets
  4. Rate of return on common stockholder’s equity
  5. Earnings per share of stock
  6. Inventory turnover
  7. Current ratio
  8. Quick ratio
  9. Accounts receivable turnover
  10. Debt to equity ratio
  11. Times interest earned
  12. What conclusions can you draw regarding this company's performance?
  13. Do you believe you have enough information to make a thorough analysis? Take into considerations the benefits of trend analysis.

(Please provide excel formulas to calculations)
Answer #1

a.profit margin ratio = (profit before taxes and interest/ Net sales)*100

=149000/420000 )*100           [140000 profit before taxes + 9000 interest]

=35.48%

b. total asset turnover ratio

net sales/ average total asset

420000/[(430000+392000)/2]

420000/411000

1.022

c.rate of return on total asset

Net Income / Average total Assets

=70000/411000

17.03%

d.Rate of return on common stockholder's equity

Net income / Average common equity

=70000/[(110000+150000)/2]

53.85%

e.eps= net income / no of shares outstanding

=70000/30000

=2.33$